OpenStudy (elise_a18):

Plutonium-240 decays according to the function Q(t)= Q sub o *e^(-kt) where Q represents the quantity remaining after t years and k is the decay constant, 0.00011... How long will it take 24 grams of plutonium-240 to decay to 20 grams?

OpenStudy (wolf1728):

Q(t)= Q sub o *e^(-kt) You'll have to solve that equation for time and that is: time = ln(Qt / Qo) / k (where "ln" mean "natural log") For one think I'd say your value of "k" should be negative: time = ln((20 / 24) / -.00011 time = ln (.83333333) / -.00011 time = -.18232/-.00011 time = 1,657 years Pages that will help you a LOT http://www.1728.org/halflif2.htm (see problem 4) http://www.1728.org/halflife.htm